The demand product is P=400-10Qd and the supply is P=200+10Qs. What is the price elasticity I’d demand for this product when the market is in equilibrium?


1
Expert's answer
2021-03-14T19:03:13-0400

Market equilibrium occurs when "P_d=P_s":


"400-10Q_E=200+10Q_E,""200=20Q_E,""Q_E=10."

Then we can find equilibrium price:


"P_E=400-10\cdot10=\$300."


We can find the elasticity of the demand when the market is in equilibrium as follows:


"E_d=-(\dfrac{1}{slope})\dfrac{P}{Q_d},""E_d=-(\dfrac{1}{-10})\dfrac{300}{10}=-3.0."